# Decomposing Global Feature Effects Based on Feature Interactions

### v2

@@ 4 Generalized Additive Decomposition of Global EffecTs (GADGET)
Our proposed framework, GADGET, additively decomposes global feature effects by minimizing feature interactions to produce regional feature effects. One or multiple features can be used to derive interpretable subregions where (aggregated) regional feature effects are more representative of individual observations within each region. Below, we will introduce the GADGET methodology, applicable to various feature effect methods, by formally defining the required axiom and demonstrating its satisfaction by popular methods.
@@ 4.1 Measuring Interaction-Related Heterogeneity
Let $h(x_{j},\mathbf{x}_{-j}^{(i)})$ be a function that quantifies the local effect of the $j$-th feature on the $i$-th observation at a specific value $x_{j}\in\mathcal{X}_{j}$, calculated by a function $h:\mathbb{R}^{p}\rightarrow\mathbb{R}$. We keep this function generic for now, but discuss in Sections 4.3 to 4.5 several specific choices, including the underlying local effect functions of PD, ALE, and SHAP dependence (SD) curves. For example, in the case of PD, $h$ is defined by the $i$-th mean-centered ICE curve of the $j$-th feature at one specific grid point $x_{j}$.
@@
$\mathcal{L}\left(\mathcal{A}_{g},x_{j}\right)$ $=$ $\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{g}}\left(h(x_{j},\mathbf{x}_{-j}^{(i)})-\mathds{E}[h(x_{j},X_{-j})|\mathcal{A}_{g}]\right)^{2}.$ (6)
The risk function $\mathcal{R}$ for the $j$-th feature and subspace $\mathcal{A}_{g}$ is defined by aggregating the point-wise losses of Eq. (6) over a given $m$-dimensional grid of feature values $\tilde{\mathbf{x}}_{j}$:
$\mathcal{R}\left(\mathcal{A}_{g},\tilde{\mathbf{x}}_{j}\right)$ $=$ $\sum\limits_{k:k\in\{1,\ldots,m\}\wedge\tilde{x}_{j}^{(k)}\in\iota_{j}}\mathcal{L}\left(\mathcal{A}_{g},\tilde{x}_{j}^{(k)}\right).$ (7)
For instance, $\tilde{\mathbf{x}}_{j}$ can be $m$ values sampled from the realizations of $j$-th feature, an equidistant grid or quantiles, but we restrict the summation in Eq. (7) to values which fall inside the considered subspace $\mathcal{A}_{g}$, so $\iota_{g,j}$.
The loss in Eq. (6) measures the heterogeneity of the local effects, which can be decomposed into the main and interaction effects of involved features. We are only interested in measuring the interaction-related heterogeneity of the feature of interest, disregarding variability caused by any other effects. This is possible when Axiom 1 holds.
@@
A local feature effect function $h:\mathbbm{R}^{p}\rightarrow\mathbbm{R}$ satisfies the local decomposability axiom if and only if the decomposition of the $i$-th local effect $h(x_{j},\mathbf{x}_{-j}^{(i)})$ of feature $\mathbf{x}_{j}$ at $x_{j}$ solely depends on main and higher-order effects of feature $\mathbf{x}_{j}$:
@@
In Sections 4.3 to 4.5 we show that this axiom is fulfilled by the most popular feature effect methods, at least when properly centered. If Axiom 1 is satisfied, then Theorem 2 guarantees that the loss function defined in Eq. (6) quantifies the interaction-related heterogeneity of local effects (feature interactions) of feature $\mathbf{x}_{j}$ at $x_{j}$ within the regarded subspace $\mathcal{A}_{g}$. Specifically, Theorem 2 shows that the main effect $g_{j}(x_{j})$ no longer affects the loss $\mathcal{L}$, rendering $\mathcal{L}$ a pure measure of interaction-related heterogeneity.
@@
If the local feature effect function $h(x_{j},\mathbf{x}_{-j}^{(i)})$ satisfies Axiom 1, then the loss function $\mathcal{L}\left(\mathcal{A}_{g},x_{j}\right)$ defined in Eq. (6) only depends on feature interactions between the feature $\mathbf{x}_{j}$ at $x_{j}$ and features in $-j$:
$\mathcal{L}\left(\mathcal{A}_{g},x_{j}\right)=\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{g}}\left(\sum\limits_{k=1}^{p-1}\;\sum\limits_{\begin{subarray}{c}W\subseteq-j,\\ |W|=k\end{subarray}}g_{W\cup j}(x_{j},\mathbf{x}_{W}^{(i)})-\mathds{E}[g_{W\cup j}(x_{j},X_{W})|\mathcal{A}_{g}]\right)^{2}.$
The proof can be found in Appendix A.1.
@@ 4.2 The GADGET Algorithm
The goal of the GADGET algorithm is to find regions with little interaction-related heterogeneity. A user specifies two sets: $S\subseteq\{1,\ldots,p\}$ containing the features of interest for which we want to receive representative regional feature effect plots by minimizing their interaction-related heterogeneity, and $Z\subseteq\{1,\ldots,p\}$ containing features assumed to interact with those in $S$. Features in $Z$ serve as split features when partitioning the feature space to minimize interactions with features in $S$. Features in $-S$ are usually assumed not to interact with other features and are thus only calculated and visualized on a global level. Algorithm 1 defines a single partitioning step of GADGET, which is inspired by the CART algorithm (Breiman et al., 1984) and is recursively repeated until a certain stop criterion is met (see Appendix C for more details). We greedily search for the best split point within the feature subset $Z$ such that the interaction-related heterogeneity of feature subset $S$ is minimized in the two resulting subspaces. The interaction-related heterogeneity is measured by the variance of the local feature effects of $\mathbf{x}_{j}$ within the new subspace (see Eq. 7). Hence, for each candidate split feature $z$ and candidate split point $t$, we sum up the risks of the two resulting subspaces for all features in $S$ (line 6 in Algorithm 1) and then choose the split $(\hat{t},\hat{z})$ that minimizes the interaction-related heterogeneity of all features $j\in S$ (line 9 in Algorithm 1). Note that Algorithm 1 is exemplarily defined for numeric features, however, GADGET can also handle categorical features which is described in detail in Appendix B.5.
1:  input: subspace $\mathcal{A}\subseteq\mathcal{X}$, risk $\mathcal{R}$, features $S$, splitting set $Z$, grids $\tilde{\mathbf{x}}_{j}\;\forall j\in S$
@@
6:        $\mathcal{I}(t,z)=\sum_{j\in S}\left(\mathcal{R}(\mathcal{A}_{l}^{t,z},\tilde{\mathbf{x}}_{j})+\mathcal{R}(\mathcal{A}_{r}^{t,z},\tilde{\mathbf{x}}_{j})\right)$
@@
55footnotetext: $S$ and $Z$ can be distinct or partially or fully overlap with each other, see also Section 5.
@@
Suppose the local feature effect function $h$ satisfies Axiom 1. If the feature subset $-Z$ does not contain any features interacting with any features indexed by $j\in S$, then after sufficiently many iterations, GADGET returns regions $\mathcal{A}_{g}$ with $\mathcal{R}(\mathcal{A}_{g},\tilde{\mathbf{x}}_{j}$) = 0 for all final regions, i.e., the theoretical minimum of the objective function $\mathcal{I}$ in Alogrithm 1 is 00. The proof can be found in Appendix A.2.
In practice, one rarely wants to run the algorithm to convergence. The question of how many partitioning steps should be performed depends on the underlying research question. If the user is more interested in reducing the interaction-related heterogeneity as much as possible, they might split rather deeply, depending on the complexity of interactions learned by the model. However, this might lead to many regions that are more challenging to interpret. If a small number of regions is of interest, one might prefer a shallow tree, thus reducing only the heterogeneity of the features that interact the most. More detailed recommendations on specifying these stop criteria are provided in Appendix C.
@@ 4.3 GADGET-PD
The PD plot is based on ICE curves (local feature effects). However, ICE curves do not satisfy Axiom 1, since the decomposition of the $i$-th ICE curve (i.e., $\hat{f}^{\mathbf{x}}_{j}(x_{j})=\hat{f}(x_{j},\mathbf{x}_{-j}^{(i)})$) also contains main or interaction effects of the $i$-th observation that are independent of feature $\mathbf{x}_{j}$ (see Appendix A.3). These effects can be canceled out by centering ICE curves w.r.t. the mean of each curve, yielding the following local effect function in GADGET-PD:
@@
Mean-centered ICE curves satisfy the local decomposibility axiom (Axiom 1). The proof is given in Appendix A.3.
@@
$\mathcal{L}^{PD}\left(\mathcal{A}_{g},x_{j}\right)=\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{g}}\left(\hat{f}^{c}(x_{j},\mathbf{x}_{-j}^{(i)})-\mathds{E}[\hat{f}^{c}(x_{j},X_{-j})|\mathcal{A}_{g}]\right)^{2}.$ (9)
In Appendix A.4, we show that the REPID method discussed in Section 3 is a special case of GADGET-PD for the loss function in Eq. (9) with only one feature of interest (i.e., $S=j$) and where we consider all other features to be potential split features (i.e., $Z=-j$).
@@
Figure 2 visualizes the global PD and ICE curves along with the regional ones after applying GADGET-PD. The global plots for features $\mathbf{x}_{1}$ and $\mathbf{x}_{3}$ show high heterogeneity in the underlying local effects, indicating that the global PD curve does not represent the respective ICE curves well. For example, the nearly horizontal global PD curve of $\mathbf{x}_{1}$ (grey line) indicates no influence on the model’s predictions. Consequently, it is not representative of the underlying local feature effects (ICE curves), which vary highly due to the feature interaction with $\mathbf{x}_{3}$. We applied GADGET-PD with $S=Z=\{1,2,3\}$ to visualize the (regional) feature effect of all available features by considering all features as possible interaction (split) features. GADGET-PD performed one split with regard to $\mathbf{x}_{3}\approx 0$. Thus, the correct split feature and its corresponding split point are found by GADGET-PD such that the interaction-related heterogeneity of all features in $S$ is almost completely reduced and only the main effects within the subspaces remain. Hence, the joint mean-centered PD function $f^{PD,c}_{S|\mathcal{A}_{g}}$ within each final subspace $\mathcal{A}_{g}$ can be decomposed into the respective 1-dimensional mean-centered PD functions as defined in Eq. (8). The decomposition is explained in more detail in Appendix B.1.
@@
Figure 2: Visualization of applying GADGET with $S=Z=\{1,2,3\}$ to mean-centered ICE curves of the stated uncorrelated simulation example with $Y=3X_{1}\mathbbm{1}_{X_{3}>0}-3X_{1}\mathbbm{1}_{X_{3}\leq 0}+X_{3}+\epsilon$ with $\epsilon\sim\mathbb{N}(0,0.09)$. The upper plots show the mean-centered ICE and PD curves on the entire feature space, while the lower plots represent the respective mean-centered ICE and PD curves after partitioning the feature space w.r.t. $\mathbf{x}_{3}=-0.003$.

### v3

@@ 4 Generalized Additive Decomposition of Global EffecTs (GADGET)
Our proposed framework, GADGET, additively decomposes global feature effects by minimizing feature interactions to produce regional feature effects. One or more features can be used to derive interpretable subregions where aggregated regional feature effects better represent individual observations within each region. Below, we will introduce the GADGET methodology, applicable to various feature effect methods, by formally defining the required axiom and demonstrating its satisfaction by popular methods.
@@ 4.1 Measuring Interaction-Related Heterogeneity
Let $h(x_{j},\mathbf{x}_{-j}^{(i)})$ be a function that quantifies the local effect of the $j$-th feature on the $i$-th observation at a specific value $x_{j}\in\mathcal{X}_{j}$, calculated by a function $h:\mathbb{R}^{p}\rightarrow\mathbb{R}$. We keep this function generic for now, but discuss in Sections 4.3 to 4.5 several specific choices, including the underlying local effect functions of PD, ALE, and SHAP dependence (SD) curves. For example, in the case of PD, $h$ is defined by the $i$-th mean-centered ICE curve of the $j$-th feature at one specific grid point $x_{j}$ (see Figure 2).
@@
$\mathcal{L}_{j}\left(\mathcal{A}_{g},x_{j}\right)$ $=$ $\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{g}}\left(h(x_{j},\mathbf{x}_{-j}^{(i)})-\mathds{E}[h(x_{j},X_{-j})|\mathcal{A}_{g}]\right)^{2}.$ (6)
Eq. (6) measures the heterogeneity of the local feature effects of feature $\mathbf{x}_{j}$ at a specific value $x_{j}$ by calculating the L2 loss as demonstrated in the left and middle plots of Figure 2. Here, the local feature effects are the mean-centered ICE curves and the respective mean-centered PD curve represents the expected effect of the feature of interest.
The risk function $\mathcal{R}_{j}$ for the $j$-th feature and subspace $\mathcal{A}_{g}$ is defined by aggregating the point-wise losses of Eq. (6) over a given $m$-dimensional grid of feature values $\tilde{\mathbf{x}}_{j}$:
$\mathcal{R}_{j}\left(\mathcal{A}_{g},\tilde{\mathbf{x}}_{j}\right)$ $=$ $\sum\limits_{k:k\in\{1,\ldots,m\}\wedge\tilde{x}_{j}^{(k)}\in\iota_{j}}\mathcal{L}_{j}\left(\mathcal{A}_{g},\tilde{x}_{j}^{(k)}\right).$ (7)
For instance, $\tilde{\mathbf{x}}_{j}$ can be $m$ values sampled from the realizations of the $j$-th feature, an equidistant grid or quantiles, but we restrict the summation in Eq. (7) to values which fall inside the considered subspace $\mathcal{A}_{g}$, so $\iota_{g,j}$. It follows that Eq. (7) summarizes the heterogeneity of local feature effects of the feature $\mathbf{x}_{j}$ for its entire feature range within the region $\mathcal{A}_{g}$ in one number and therefore quantifies how much the feature’s influence of individual observations varies from the expected (average) feature’s influence on the predictions (e.g., right plot of Figure 2 in the case of mean-centered ICE and PD curves).
The loss in Eq. (6) measures the heterogeneity of the local effects, which can be decomposed into the main and interaction effects of involved features (see Eq. (1)). Since we are only interested in measuring the interaction-related heterogeneity of the feature of interest, disregarding variability caused by any other effects, the local feature effect function $h$ needs to be defined in such a way that the heterogeneity of local feature effects measured by Eq. (6) and  (7) are solely based on feature interactions between the feature of interest $\mathbf{x}_{j}$ and the remaining feature in the data set. This is possible when Axiom 1 holds. [^6] We show in Sections 4.3 to 4.5 that common feature effect methods satisfy this general decomposition based on the functional ANOVA decomposition in Eq. (1).
@@
A local feature effect function $h:\mathbbm{R}^{p}\rightarrow\mathbbm{R}$ satisfies the local decomposability axiom if and only if $\forall i\in\{1,\dots,n\}$ the decomposition of the $i$-th local effect $h(x_{j},\mathbf{x}_{-j}^{(i)})$ of feature $\mathbf{x}_{j}$ at $x_{j}$ solely depends on main and higher-order effects of feature $\mathbf{x}_{j}$:
@@
In Sections 4.3 to 4.5 we show that this axiom is fulfilled by the most popular feature effect methods, at least when properly centered. If Axiom 1 is satisfied, then Theorem 2 guarantees that the loss function defined in Eq. (6) quantifies the interaction-related heterogeneity of local effects (feature interactions) of feature $\mathbf{x}_{j}$ at $x_{j}$ within the regarded subspace $\mathcal{A}_{g}$. Specifically, Theorem 2 shows that the main effect $g_{j}(x_{j})$ no longer affects the loss $\mathcal{L}_{j}$, rendering $\mathcal{L}_{j}$ a pure measure of interaction-related heterogeneity.
@@
If the local feature effect function $h(x_{j},\mathbf{x}_{-j}^{(i)})$ satisfies Axiom 1, then the loss function $\mathcal{L}_{j}\left(\mathcal{A}_{g},x_{j}\right)$ defined in Eq. (6) only depends on feature interactions between the feature $\mathbf{x}_{j}$ at $x_{j}$ and features in $-j$:
$\mathcal{L}_{j}\left(\mathcal{A}_{g},x_{j}\right)=\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{g}}\left(\sum\limits_{k=1}^{p-1}\;\sum\limits_{\begin{subarray}{c}W\subseteq-j,\\ |W|=k\end{subarray}}g_{W\cup j}(x_{j},\mathbf{x}_{W}^{(i)})-\mathds{E}[g_{W\cup j}(x_{j},X_{W})|\mathcal{A}_{g}]\right)^{2}.$
The proof can be found in Appendix B.1.
Illustration.
To better understand Axiom 1, Theorem  2, Eq. 6 and  7, we provide an illustrative example. Let $X_{1},X_{2},X_{3}\sim\mathbbm{U}(-1,1)$ be independent, with the true underlying relationship defined by $Y=X_{1}+X_{2}+X_{3}+X_{1}X_{2}+\epsilon$, where $\epsilon\sim\mathbb{N}(0,0.0025)$. We draw 20 observations and fit a correctly specified linear model to the data. We define the local feature effect function as the mean-centered ICE curves (Section 4.3) and use $\mathbf{x}_{1}$ as the feature of interest. The $i$-th ICE curve value at grid point $\tilde{x}_{1}$ is given by $\hat{f}(\tilde{x}_{1},\mathbf{x}_{-1}^{(i)})=\tilde{x}_{1}+x_{2}^{(i)}+x_{3}^{(i)}+\tilde{x}_{1}x_{2}^{(i)}$, which corresponds to the decomposition in Eq. (1). Since this decomposition includes effects independent of feature $\mathbf{x}_{1}$ (i.e., $g_{3}(x_{3}^{(i)})=x_{3}^{(i)}$), Axiom 1 does not hold. Therefore, the ICE curve needs to be centered by its mean: $\hat{f}^{c}(\tilde{x}_{1},\mathbf{x}_{-1}^{(i)})=\hat{f}(\tilde{x}_{1},\mathbf{x}_{-1}^{(i)})-\mathds{E}[\hat{f}(X_{1},\mathbf{x}_{-1}^{(i)}]=\tilde{x}_{1}-\mathds{E}[X_{1}]+(\tilde{x}_{1}-\mathds{E}[X_{1}])x_{2}^{(i)}\approx\tilde{x}_{1}+\tilde{x}_{1}x_{2}^{(i)}$ which satisfies Axiom 1 with main effect $g_{j}(x_{j})=\tilde{x}_{1}$ and the respective higher-order effect $g_{W\cup j}(x_{j},\mathbf{x}_{W}^{(i)})=\tilde{x}_{1}x_{2}^{(i)}$. Figure 2 shows the mean-centered ICE curves and the mean-centered PD curve, which is calculated as the average over all mean-centered ICE curves. The loss function in Eq. (6) aggregates the squared distance between each mean-centered ICE curve and the PD curve at a specific grid point as demonstrated in the left and middle plot of Figure 2 for $\tilde{x}_{1}=-1$. Since mean-centered ICE curves satisfy Axiom 1, this distance only measures interaction-related heterogeneity as stated in Theorem 2: $h(x_{1},\mathbf{x}_{-1}^{(i)})-\mathds{E}[h(x_{1},X_{-1})|\mathcal{X}]=\tilde{x}_{1}+\tilde{x}_{1}x_{2}^{(i)}-\tilde{x}_{1}-\tilde{x}_{1}\mathds{E}[X_{2}]=\tilde{x}_{1}x_{2}^{(i)}-\tilde{x}_{1}\mathds{E}[X_{2}]\approx\tilde{x}_{1}x_{2}^{(i)}$. Thus, the loss function and therefore also the risk function in Eq. (7), which aggregates over all valid grid points of the feature $\mathbf{x}_{j}$ within the respective region (e.g., right plot Figure 2), solely depend on the interaction effect between the feature of interest ($\mathbf{x}_{1}$) and other features in the data set (here: $\mathbf{x}_{2}$).
Figure 2: Mean-centered ICE and PD curves of feature $\mathbf{x}_{1}$ of the described simulation example. Left: Illustrates the distance $d^{(1)}$ that is calculated within Eq. (6) between the local feature effect (here: $h(x_{j},\mathbf{x}_{-j}^{(i)}):=\hat{f}^{c}(x_{j},\mathbf{x}_{-j}^{(i)})$, i.e., the mean-centered ICE curve) and the expected effect within region $\mathcal{A}_{g}$ (here: $\mathds{E}[h(x_{j},X_{-j})|\mathcal{X}]=\mathds{E}[\hat{f}^{c}(x_{j},X_{-j})|\mathcal{X}]$, i.e. the global mean-centered PD) at the first grid point $\tilde{x}_{1}=-1$ for the first observation. Middle: The distances $d$ are calculated for all mean-centered ICE curves at the first grid point and the squared values are summed up to obtain the loss function value for the first grid point as defined in Eq. (6). Right: The risk function value is calculated according to Eq. (7) by aggregating the loss function values over the valid grid points. This measures the heterogeneity of the mean-centered ICE curves of feature $\mathbf{x}_{1}$, which quantifies the interaction-related heterogeneity between $\mathbf{x}_{1}$ and $\mathbf{x}_{-1}$ based on Theorem 2.
@@ 4.2 The GADGET Algorithm
The goal of the GADGET algorithm is to find regions with little interaction-related heterogeneity. A user specifies two sets: $S\subseteq\{1,\ldots,p\}$ containing the features of interest for which we want to receive representative regional feature effect plots by minimizing their interaction-related heterogeneity, and $Z\subseteq\{1,\ldots,p\}$ containing features assumed to interact with those in $S$. Features in $Z$ serve as split features when partitioning the feature space to minimize interactions with features in $S$. Features in $-S$ are usually assumed not to interact with other features and are thus only calculated and visualized on a global level. Algorithm 1 defines a single partitioning step of GADGET, which is inspired by the CART algorithm (Breiman et al., 1984) and is recursively repeated until a certain stop criterion is met (see Appendix D for more details). We greedily search for the best split point within the feature subset $Z$ such that the interaction-related heterogeneity of feature subset $S$ is minimized in the two resulting subspaces. The interaction-related heterogeneity is measured by the variance of the local feature effects of $\mathbf{x}_{j}$ within the new subspace (see Eq. 7). Hence, for each candidate split feature $z$ and candidate split point $t$, we sum up the risks of the two resulting subspaces for all features in $S$ (line 6 in Algorithm 1) and then choose the split $(\hat{t},\hat{z})$ that minimizes the interaction-related heterogeneity of all features $j\in S$ (line 9 in Algorithm 1). Note that Algorithm 1 is exemplarily defined for numeric features, however, GADGET can also handle categorical features which is described in detail in Appendix C.5.
1:  input: subspace $\mathcal{A}\subseteq\mathcal{X}$, risk $\mathcal{R}_{j}$, features $S$, splitting set $Z$, grids $\tilde{\mathbf{x}}_{j}\;\forall j\in S$
@@
6:        $\mathcal{I}(t,z)=\sum_{j\in S}\left(\mathcal{R}_{j}(\mathcal{A}_{l}^{t,z},\tilde{\mathbf{x}}_{j})+\mathcal{R}_{j}(\mathcal{A}_{r}^{t,z},\tilde{\mathbf{x}}_{j})\right)$
@@
66footnotetext: $S$ and $Z$ can be distinct or partially or fully overlap with each other, see also Section 5.
@@
Suppose the local feature effect function $h$ satisfies Axiom 1. If the feature subset $-Z$ does not contain any features interacting with any features indexed by $j\in S$, i.e., $\forall j\in S$ and $\forall L\subseteq-Z$ with $-Z\subset\{1,\ldots,p\}\setminus j$ and $\forall K\subseteq Z:=\{\emptyset,1,\ldots,p\}\setminus\{-Z\cup j\}$ it holds: $g_{L\cup K\cup j}(X_{j},X_{L},X_{K})=0$, then after sufficiently many iterations, GADGET returns regions with $\mathcal{R}_{j}(\mathcal{A}_{g},\tilde{\mathbf{x}}_{j}$) = 0 for every final region $\mathcal{A}_{g}$, i.e., the theoretical minimum of the objective function $\mathcal{I}$ in Algorithm 1 is 00. The proof can be found in Appendix B.2.
In practice, one rarely wants to run the algorithm to convergence. The question of how many partitioning steps should be performed depends on the underlying research question. If the user is more interested in reducing the interaction-related heterogeneity as much as possible, they might split rather deeply, depending on the complexity of interactions learned by the model. However, this might lead to many regions that are more challenging to interpret. If a small number of regions is of interest, one might prefer a shallow tree, thus reducing only the heterogeneity of the features that interact the most. More detailed recommendations on specifying these stop criteria are provided in Appendix D.
@@ 4.3 GADGET-PD
The PD plot is based on ICE curves (local feature effects). However, ICE curves do not satisfy Axiom 1, since the decomposition of the $i$-th ICE curve (i.e., $\hat{f}^{\mathbf{x}}_{j}(x_{j})=\hat{f}(x_{j},\mathbf{x}_{-j}^{(i)})$) also contains main or interaction effects of the $i$-th observation that are independent of feature $\mathbf{x}_{j}$ (see Appendix B.3). These effects can be canceled out by centering ICE curves w.r.t. the mean of each curve, yielding the following local effect function in GADGET-PD:
@@
Mean-centered ICE curves satisfy the local decomposability axiom (Axiom 1). The proof is given in Appendix B.3.
@@
$\mathcal{L}_{j}^{PD}\left(\mathcal{A}_{g},x_{j}\right)=\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{g}}\left(\hat{f}^{c}(x_{j},\mathbf{x}_{-j}^{(i)})-\mathds{E}[\hat{f}^{c}(x_{j},X_{-j})|\mathcal{A}_{g}]\right)^{2}.$ (9)
In Appendix B.4, we show that the REPID method discussed in Section 3 is a special case of GADGET-PD for the loss function in Eq. (9) with only one feature of interest (i.e., $S=j$) and where we consider all other features to be potential split features (i.e., $Z=-j$).
@@
Figure 3 visualizes the global PD and ICE curves along with the regional ones after applying GADGET-PD. The global plots for features $\mathbf{x}_{1}$ and $\mathbf{x}_{3}$ show high heterogeneity in the underlying local effects, indicating that the global PD curve does not represent the respective ICE curves well. For example, the nearly horizontal global PD curve of $\mathbf{x}_{1}$ (grey line) indicates no influence on the model’s predictions. Consequently, it is not representative of the underlying local feature effects (ICE curves), which vary highly due to the feature interaction with $\mathbf{x}_{3}$. We applied GADGET-PD with $S=Z=\{1,2,3\}$ to visualize the (regional) feature effect of all available features by considering all features as possible interaction (split) features. GADGET-PD performed one split with regard to $\mathbf{x}_{3}\approx 0$. Thus, the correct split feature and its corresponding split point are found by GADGET-PD such that the interaction-related heterogeneity of all features in $S$ is almost completely reduced and only the main effects within the subspaces remain. Hence, the joint mean-centered PD function $f^{PD,c}_{S|\mathcal{A}_{g}}$ within each final subspace $\mathcal{A}_{g}$ can be decomposed into the respective 1-dimensional mean-centered PD functions as defined in Eq. (8). The decomposition is explained in more detail in Appendix C.1.
@@
Figure 3: Visualization of applying GADGET with $S=Z=\{1,2,3\}$ to mean-centered ICE curves of the uncorrelated simulation example with $Y=3X_{1}\mathbbm{1}_{X_{3}>0}-3X_{1}\mathbbm{1}_{X_{3}\leq 0}+X_{3}+\epsilon$ where $\epsilon\sim\mathbb{N}(0,0.09)$. Plots show mean-centered ICE and PD curves on the entire feature space (upper) and within regions after partitioning the feature space w.r.t. $\mathbf{x}_{3}=-0.003$ (lower).